OpenStudy (anonymous):

find polar coordinates for three different points that line on the graph of the equation r=-4cosθ. I have (-4,0) but i can't find the other two.

OpenStudy (rational):

plugin \(\theta = \dfrac{\pi}{2}\), you get \(r = -4\cos(\dfrac{\pi}{2}) = -4*0 = 0\) so (0, pi/2) is another point on the given graph

OpenStudy (rational):

similarly see if you can find another point by plugging in \(\large \theta = \dfrac{\pi}{6}\)

OpenStudy (anonymous):

almost -4 (3.999) so the next point would be (-4, pi/6)?

OpenStudy (rational):

doesn't look correct

OpenStudy (rational):

\(r = -4\cos(\theta)\) plugin \(\theta = \dfrac{\pi}{6}\) \(r = -4\cos(\dfrac{\pi}{6}) = -4*\dfrac{1}{2} = -2\) so (-2, pi/6) is another point on the graph

OpenStudy (anonymous):

had the calculator in degree mode...

OpenStudy (anonymous):

you're using the unit circle aren't you?

OpenStudy (anonymous):

cos(pi/3)= 1/2...

OpenStudy (anonymous):

thank you

OpenStudy (rational):

thats it!